# PDE-aware Optimizer for Physics-informed Neural Networks

### v1

@@ 1 Introduction
A significant advancement in scientific machine learning is physics-informed machine learning, which integrates physical laws and domain-specific constraints directly into the learning process. This integration can be achieved in various ways, including modifications to key components such as the model architecture, loss functions, optimisation algorithms, and hyperparameter tuning [Karniadakis2021PhysicsInformedML, Raissi2019PINNs]. A particularly effective strategy within physics-informed machine learning is embedding physical principles such as conservation laws, symmetries, invariances, and equivariances into the machine learning framework. Among the most flexible and widely adopted approaches is the use of tailored loss functions. These loss functions encode the physical constraints as soft penalties, guiding the model to learn solutions that are not only data-driven but also consistent with known physical laws. This approach has led to the development of Physics-Informed Neural Networks (PINNs), which have been successfully applied to both forward and inverse problems involving partial differential equations (PDE).
While PINNs offer conceptual simplicity and implementation flexibility, training them effectively for stiff or ill-conditioned PDEs remains challenging. Traditional first-order optimization methods, such as Adam, often struggle with convergence or require extensive tuning. To address these issues, researchers have turned to second-order optimization techniques, which utilize curvature information to make more informed updates to the model parameters [Wang2025GradientAlignment, Wang2021GradientPathologies]. Second-order optimizers such as Second-Order Adaptive Optimisation (SOAP) and Second-order Clipping Shampoo (SHAMPOO) have demonstrated superior convergence properties in training PINNs, particularly on complex PDE problems [Wang2025GradientAlignment]. These methods leverage second-order information, such as approximations of the Hessian or the Fisher information matrix, to scale the gradient updates more precisely along each parameter direction, resulting in improved training stability and accuracy. However, these benefits come at a significant cost.
The primary drawback of second-order methods lies in their high computational and memory overhead. Maintaining and updating curvature matrices—often of size proportional to the number of parameters squared—is resource-intensive, especially for deep networks or large-scale problems [Martens2010HessianFree, Martens2015KFAC]. Moreover, operations such as matrix inversion or eigendecomposition add to the computational complexity [RoostaKhorasani2016SubsampledNewton]. Despite these challenges, ongoing research aims to make second-order methods more tractable through matrix sketching, low-rank, and block-diagonal approximations [Gupta2018Shampoo, Yao2021AdaHessian].
@@ 3 Gradient Misalignment in PINNs
The two distinct types of gradient conflicts [Wang2025GradientAlignment]:
@@ 4 PDE-Aware Optimizer
3:$\bm{m}_{-1}\leftarrow\bm{0}$triangleright first moment
4:$\bm{v}_{-1}\leftarrow\bm{0}$triangleright second moment
5:for $t=0$ to $T-1$ do
@@ 5 Experimental Setup
Batch size $B=1{,}024$; epochs $T=10{,}000$.
@@
Hyper-parameter tuning:
We carried out a full $2\times 2\times 2$ grid search over the learning-rate $\eta\!\in\!\{10^{-3},10^{-4}\}$ and moments $\beta_{1}\!\in\!\{0.9,0.99\}$, $\beta_{2}\!\in\!\{0.99,0.999\}$ for our PDE-aware optimizer on the 1-D Burgers equation. We chose to begin at the default values of the Adam optimizer and vary the hyperparameters from there. The rationale behind the choice for $\beta_{2}$ was to increase $1-\beta_{2}$ since that term would determine how much the corresponding residual term would be weighted in $v_{t}$. Figure 3 visualizes the validation loss surface for the eight configurations. The best setting—$\eta=10^{-3},\;\beta_{1}=0.99,\;\beta_{2}=0.99$—achieved an average validation loss of $4.32\times 10^{-2}$, with the two closest neighbors ($\eta=10^{-3},\beta_{1}=0.99,\beta_{2}=0.999$ and $\eta=10^{-3},\beta_{1}=0.9,\beta_{2}=0.99$) trailing by less than $3\times 10^{-4}$.

### v2

@@ 1 Introduction
A significant advancement in scientific machine learning is physics-informed machine learning, which integrates physical laws and domain-specific constraints directly into the learning process. This integration can be achieved in various ways, including modifications to key components such as the model architecture, loss functions, optimisation algorithms, and hyperparameter tuning [undefa, undefd]. A particularly effective strategy within physics-informed machine learning is embedding physical principles such as conservation laws, symmetries, invariances, and equivariances into the machine learning framework. Among the most flexible and widely adopted approaches is the use of tailored loss functions. These loss functions encode the physical constraints as soft penalties, guiding the model to learn solutions that are not only data-driven but also consistent with known physical laws. This approach has led to the development of Physics-Informed Neural Networks (PINNs), which have been successfully applied to both forward and inverse problems involving partial differential equations (PDE).
While PINNs offer conceptual simplicity and implementation flexibility, training them effectively for stiff or ill-conditioned PDEs remains challenging. Traditional first-order optimization methods, such as Adam, often struggle with convergence or require extensive tuning. To address these issues, researchers have turned to second-order optimization techniques, which utilize curvature information to make more informed updates to the model parameters [undefg, undeff]. Second-order optimizers such as Second-Order Adaptive Optimisation (SOAP) and Second-order Clipping Shampoo (SHAMPOO) have demonstrated superior convergence properties in training PINNs, particularly on complex PDE problems [undefg]. These methods leverage second-order information, such as approximations of the Hessian or the Fisher information matrix, to scale the gradient updates more precisely along each parameter direction, resulting in improved training stability and accuracy. However, these benefits come at a significant cost.
The primary drawback of second-order methods lies in their high computational and memory overhead. Maintaining and updating curvature matrices—often of size proportional to the number of parameters squared—is resource-intensive, especially for deep networks or large-scale problems [undefb, undefc]. Moreover, operations such as matrix inversion or eigendecomposition add to the computational complexity [undefe]. Despite these challenges, ongoing research aims to make second-order methods more tractable through matrix sketching, low-rank, and block-diagonal approximations [undef, undefh].
@@ 3 Gradient Misalignment in PINNs
The two distinct types of gradient conflicts [undefg]:
@@ 4 PDE-Aware Optimizer
[…]
5:for $t=0$ to T1T-1 do
@@ 5 Experimental Setup
Batch size B=1,024B=1{,}024; epochs T=10, [^0] T=10{,}000.
@@
$\text{Rel-}L^{2}=\Bigl(\tfrac{\sum(u_{\theta}-u_{\text{exact}})^{2}}{\sum u_{\text{exact}}^{2}}\Bigr)^{1/2}.$
Hyper-parameter tuning: